OpenStudy (anonymous):

y + 3 = 2x y + 4 = 3x If you use the comparison method to solve the given system, what is the resulting equation? y + 3 = y + 4 2x - 3 = 3x - 4 2x + 3 = 3x + 4

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

y + 3 = y + 4?

OpenStudy (trojanpoem):

y + 3 + 1 = 3x 2x + 1 = 3x x = 1 2 + 3 = 3 + 4 (false) 2- 3 = 3 - 4 (true) 2x - 3 = 3x - 4

OpenStudy (anonymous):

thanks

OpenStudy (trojanpoem):

yw.